evaluate the following integral
∫ e^(3x) / e^(3x) + 1 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{e^{3x}}{e^{3x}+1}dx
e^{3x}+1 = t
3e^{3x}dx = dt
I = \frac{1}{3}\int \frac{1}{t}dt
I = \frac{1}{3}lnt + c
I = \frac{1}{3}ln(e^{3x}+1) + c